### v3

# A Survey of Test-Time Compute: From Intuitive Inference to Deliberate Reasoning

Abstract.
The remarkable performance of the o1 model in complex reasoning demonstrates that test-time compute scaling can further unlock the model’s potential, enabling powerful System-2 thinking. However, there is still a lack of comprehensive surveys for test-time compute scaling. We trace the concept of test-time compute back to System-1 models. In System-1 models, test-time compute addresses distribution shifts and improves robustness and generalization through parameter updating, input modification, representation editing, and output calibration. In System-2 models, it enhances the model’s reasoning ability to solve complex problems through repeated sampling, self-correction, and tree search. We organize this survey according to the trend of System-1 to System-2 thinking, highlighting the key role of test-time compute in the transition from System-1 models to weak System-2 models, and then to strong System-2 models. We also point out advanced topics and future directions. [^1] https://github.com/Dereck0602/Awesome_Test_Time_LLMs.
Large Language Models, Test-time compute, Reasoning
ccs: Computer methodologies Artificial intelligence; Nature language processing; Nature language generation
1. Introduction
Over the past decades, deep learning with its scaling effects has been the driving engine behind the artificial intelligence revolution. Particularly in the text modality, large language models (LLMs) represented by the GPT series (Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2023) have demonstrated that larger models and more training data lead to better performance on downstream tasks. However, on the one hand, further scaling in the training phase becomes difficult due to the scarcity of data and computational resources (Villalobos et al., 2024); on the other hand, existing models still perform far below expectations in terms of robustness and handling complex tasks. These shortcomings are attributed to the model’s reliance on fast, intuitive System-1 thinking, rather than slow, deep System-2 thinking (Weston and Sukhbaatar, 2023). Recently, large reasoning models (LRMs), represented by OpenAI-o1/o3 (OpenAI, 2024), DeepSeek-R1 (Guo et al., 2025a), and Gemini 2.5 (DeepMind, 2025), equipped with System-2 thinking, have gained attention for their outstanding performance in complex reasoning tasks. It demonstrates a test-time compute scaling effect: the greater the computational effort in the inference, the better the model’s performance.
The concept of test-time compute emerged before the rise of LLMs and was initially applied to System-1 models (illustrated in Figure 1). These System-1 models can only perform limited perceptual tasks, relying on patterns learned during training for predictions. As a result, they are constrained by the assumption that training and testing are identically distributed and lack robustness and generalization to distribution shifts (Zhuang et al., 2020). Many works have explored test-time adaptation (TTA) to improve model robustness by updating parameters (Wang et al., 2021; Ye et al., 2023), modifying the input (Dong et al., 2024a), editing representations (Rimsky et al., 2024), and calibrating the output (Zhang et al., 2023c). With TTA, the System-1 model slows down its thinking process and has better generalization. However, TTA is an implicit slow thinking, unable to exhibit explicit, logical thinking process like humans, and struggles to handle complex reasoning tasks. Thus, TTA-enabled models perform weak System-2 thinking.
Figure 1. Illustration of test-time compute in the System-1 and System-2 model.
Currently, advanced LLMs with chain-of-thought (CoT) prompting (Wei et al., 2022) have enabled language models to perform explicit System-2 thinking (Hagendorff et al., 2023). However, vanilla CoT is limited by error accumulation and linear thinking pattern (Stechly et al., 2024), making it difficult to fully simulate non-linear human cognitive processes such as brainstorming, reflection, and backtracking. To achieve stronger System-2 models, researchers employ test-time compute strategies to extend model reasoning’s breadth, depth and accuracy, such as repeated sampling (Cobbe et al., 2021), self-correction (Shinn et al., 2023), and tree search (Yao et al., 2023). Repeated sampling simulates the diversity of human thinking, self-correction enables LLMs to reflect, and tree search enhances reasoning depth and backtracking.
To the best of our knowledge, this paper is the first to systematically review test-time compute methods and thoroughly explore their critical role in advancing models from System-1 to weak System-2, and ultimately to strong System-2 thinking. In Section 2, we present the background of System-1 and System-2 thinking. Section 3 and Section 4 detail the test-time compute methods for the System-1 and System-2 models. Then, we discuss advanced topics and future directions in Section 5 and 6. Additionally, we review benchmarks in Section 7.
2. Background
{forest}
forked edges, for tree= child anchor=west, parent anchor=east, grow’=east, anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7em,font=,, where level=2text width=7em,font=,, where level=3text width=10em,font=,, where level=4text width=12em,font=,, [ Test-time Computing, ver [  Test-time
 Adaptation (§3) [  Parameter
 Updating [ Test-time Training [  TTT (Sun et al., 2020), TTT++ (Liu et al., 2021), CPT (Zhu et al., 2024); etc. , leaf, text width=34em ] ] [ Fully TTA [  Tent (Wang et al., 2021), SAR (Niu et al., 2023), TPT (Shu et al., 2022), OIL (Ye et al., 2022), Anti-CF (Su et al., 2023), RLCF (Zhao et al., 2024a); etc. , leaf, text width=34em ] ] ] [  Input
 Modification [ Demonstration Selection [  EPR (Rubin et al., 2022), UDR (Li et al., 2023b), CQL (Zhang et al., 2022a), Entropy (Lu et al., 2022), MDL (Wu et al., 2023), HiAR (Wu et al., 2024a); etc. , leaf, text width=34em ] ] [ Demonstration Creation [  SG-ICL (Kim et al., 2022), Self-ICL (Chen et al., 2023), DAIL (Su et al., 2024), Auto-CoT (Zhang et al., 2023d), DAWN-ICL (Tang et al., 2024); etc. , leaf, text width=34em ] ] ] [  Representation
 Editing [  ITI (Li et al., 2023c), ActAdd (Turner et al., 2024), SEA (Qiu et al., 2024b), CAA (Rimsky et al., 2024); etc. , leaf, text width=45.6em ] ] [  Output
 Calibration [  $k$NN-MT (Khandelwal et al., 2021), AdaNPC (Zhang et al., 2023c), Bi-$k$NN (You et al., 2024); etc. , leaf, text width=45.6em ] ] ] [  Test-time
 Reasoning (§4) [  Feedback
 Modeling [ Score-based [  Bradley and Terry (1952), ORM (Cobbe et al., 2021), PAV (Setlur et al., 2024), PRM (Lightman et al., 2024), OmegaPRM (Luo et al., 2024); etc. , leaf, text width=34em ] ] [ Generative-based [  LLM-as-a-Judge (Zheng et al., 2023), Auto-J (Li et al., 2024f), Prometheus (Kim et al., 2024c, d), Fennec (Liang et al., 2024b),
 GenRM (Zhang et al., 2024c), CriticRM (Yu et al., 2024a); etc. , leaf, text width=34em ] ] ] [  Search
 Strategies [ Repeated Sampling [  SC-CoT (Wang et al., 2023b), PROVE (Toh et al., 2024)Cobbe et al. (2021), DiVeRSe (Li et al., 2023a), PRS (Ye and Ng, 2024); etc. , leaf, text width=34em ] [  Improvement training: ReST (Gulcehre et al., 2023), vBoN (Amini et al., 2024),  BoNBoN(Gui et al., 2024),  Chow et al. (2024); etc. , leaf, text width=34em ] ] [ Self-correction [  Self-debug (Chen et al., 2024c), RIC (Kim et al., 2023), Critic (Gou et al., 2024), Shepherd (Wang et al., 2023c), MAD (Liang et al., 2024a), IoE (Li et al., 2024a),
 Refiner (Paul et al., 2024), Reflexion (Shinn et al., 2023), Du et al. (2023b), ProgCo (Song et al., 2025b), S1 (Muennighoff et al., 2025); etc. , leaf, text width=34em ] [  Improvement training: GLoRe (Havrilla et al., 2024), SCoRe(Kumar et al., 2024), Self-correct (Welleck et al., 2023),
 Qu et al. (2024)Zhang et al. (2024a), GRPO (Shao et al., 2024), T1 (Hou et al., 2025); etc. , leaf, text width=34em ] ] [ Tree Search [  ToT (Yao et al., 2023), RAP (Hao et al., 2023), rStar (Qi et al., 2024a), TS-LLM (Feng et al., 2024a), AlphaMATH (Chen et al., 2024b); etc. , leaf, text width=34em ] [  Improvement training: ReST-MCTS* (Zhang et al., 2024m), Qin et al. (2024a),
 MCTS-DPO (Xie et al., 2024), Zhao et al. (2024b)Zhang et al. (2024h); etc. , leaf, text width=34em ] ] ] ] [  Future
 Directions (§6) [  Generalization [  Jia (2024), GRM (Yang et al., 2024a), DogeRM (Lin et al., 2024c), GenPRM (Zhao et al., 2025a), RM-R1 (Chen et al., 2025b), Search-o1 (Li et al., 2025b), Deep Research (Li et al., 2025e); etc. , leaf, text width=45.6em ] ] [  Multi-modal [  MM-CoT (Zhang et al., 2024j), VoT (Wu et al., 2024b), Lee et al. (2024), LLaVA-CoT (Xu et al., 2024a), VisualPRM (Wang et al., 2025a), LMM-R1 (Peng et al., 2025b); etc. , leaf, text width=45.6em ] ] [  Efficient [  Damani et al. (2024), OSCA (Zhang et al., 2024l), Wang et al. (2024b), CCoT (Cheng and Durme, 2024), DEER (Yang et al., 2025b), L1 (Aggarwal and Welleck, 2025), O1-Pruner (Luo et al., 2025c); etc. , leaf, text width=45.6em ] ] [  Scaling Law [  Brown et al. (2024), Snell et al. (2024), Wu et al. (2024c), Chen et al. (2024d); etc. , leaf, text width=45.6em ] ] [  Combination [  Marco-o1 (Zhao et al., 2024b), TTT (Akyürek et al., 2024), HiAR-ICL (Wu et al., 2024a); etc. , leaf, text width=45.6em ] ] ] ]
Figure 2. Taxonomy of test-time computing methods and future directions.
System-1 and System-2 thinking are psychological concepts (Kahneman, 2011). When recognizing familiar patterns or handling simple problems, humans often respond intuitively. This automatic, fast thinking is called System-1 thinking. In contrast, when dealing with complex problems like mathematical proofs or logical reasoning, deep and deliberate thought is required, referred as System-2 thinking—slow and reflective. In the field of artificial intelligence, researchers also use these terms to describe different types of models (LeCun, 2022). System-1 models respond directly based on internally encoded perceptual information and world knowledge without showing any intermediate decision-making process. In contrast, System-2 models explicitly generate reasoning processes and solve tasks incrementally. Before the rise of LLMs, System-1 models were the mainstream in AI. Although many deep learning models, such as ResNet, Transformer, and BERT, achieve excellent performance in various tasks in computer vision and natural language processing, these System-1 models, similar to human intuition, lack sufficient robustness and are prone to errors (Geirhos et al., 2020; Wang et al., 2022c; Du et al., 2023a). Nowadays, the strong generation and reasoning capabilities of LLMs make it possible to build System-2 models. Wei et al. (2022) propose the CoT, which allows LLMs to generate intermediate reasoning steps progressively during inference. Empirical and theoretical results show that this approach significantly outperforms methods that generate answers directly (Kojima et al., 2022; Zhou et al., 2023; Feng et al., 2024b; Li et al., 2024e). However, current System-2 models represented by CoT prompting still have shortcomings. The intermediate processes generated by LLMs may contain errors, leading to cumulative mistakes and ultimately resulting in incorrect answers. Although retrieval-augmented generation (RAG) helps mitigate factual errors (Trivedi et al., 2023; Guan et al., 2024; Wang et al., 2024h), their impact on improving reasoning abilities remains limited. As a result, CoT-enabled LLMs are still at the weak system-2 thinking stage.
3. Test-time Adaptation for System-1 Thinking
Figure 3. Illustration of various kinds of test-time adaptation methods.
3.1. Updating the Model
Model updating utilizes test sample information to finetune model parameters during the inference stage, enabling the model to adapt to the test distribution. The key lies in obtaining test distribution information to provide learning signals and using appropriate parameters and optimization algorithms to achieve efficient and stable updates.
Learning signal
In the inference stage, the ground-truth of test samples is unavailable. Thus, many works attempt to design unsupervised or self-supervised objectives as learning signals. Existing learning signals can be classified into two categories based on whether the training process can be modified: test-time training (TTT) and fully test-time adaptation (FTTA). TTT assumes users can modify the training process by incorporating distribution-shift-aware auxiliary tasks. During test-time adaptation, the auxiliary task loss serves as the learning signal for optimization. Many self-supervised tasks have been shown to be effective as auxiliary tasks in image modality, such as rotation prediction (Sun et al., 2020), meta learning (Bartler et al., 2022), masked autoencoding (Gandelsman et al., 2022) and contrastive learning  (Liu et al., 2021; Chen et al., 2022). Among them, contrast learning has been successfully applied to test-time adaptation for visual-language tasks due to its generalization of self-supervised learning within and across modalities (Zhu et al., 2024).
In contrast, FTTA is free from accessing the training process and instead uses internal or external feedback on test samples as learning signals. Uncertainty is the most commonly learned signal, driven by the motivation that when test samples shift from the training distribution, the model’s confidence in its predictions is lower, resulting in higher uncertainty. Tent (Wang et al., 2021) uses the entropy of model predictions as a measure of uncertainty and updates the model by minimizing the entropy. MEMO (Zhang et al., 2022b) augments the data for a single test sample and then minimizes its marginal entropy, which is more stable compared to Tent in the single-sample TTA setting. However, minimizing entropy also has pitfalls, as blindly reducing prediction uncertainty may cause the model to collapse and make trivial predictions (Press et al., 2024; Zhao et al., 2023; Su et al., 2023). Some works propose new regularization terms for minimizing entropy to avoid model collapse, including Kullback-Leibler divergence (Su et al., 2023), moment matching (Hassan et al., 2023) and entropy matching (Bar et al., 2024). For specific tasks, a small amount of human feedback or external model rewards can also serve as high-quality learning signals. Gao et al. (2022) and Li et al. (2022b) utilize user feedback to adapt the QA model. Zhan et al. (2023) apply test-time adaptation to multilingual machine translation tasks by using COMET (Rei et al., 2020) for evaluating translation quality. In cross-modal tasks such as image-text retrieval and image captioning, RLCF (Zhao et al., 2024a) demonstrates its effectiveness by using CLIP scores (Radford et al., 2021) as TTA signals. In language modeling, training with relevant contextual text at test time can reduce perplexity (Hardt and Sun, 2024; Wang et al., 2024i). Hübotter et al. (2025) theoretically shows that it reduces the uncertainty of test samples and proposes a more effective active learning selection strategy.
Updating parameters
To advance the application of TTA in real-world scenarios, researchers must address challenges of efficiency and stability. To improve efficiency, many methods only fine-tune a small subset of parameters, such as normalization layers (Schneider et al., 2020), soft prompt (Lester et al., 2021; Shu et al., 2022; Hassan et al., 2023; MA et al., 2023; Feng et al., 2023; Niu et al., 2024), low-rank module (Hu et al., 2022; Imam et al., 2024; Hu et al., 2025), adapter module (Houlsby et al., 2019; Muhtar et al., 2024; Su et al., 2023) and cross-modality projector (Zhao et al., 2024a). Although the number of parameters to fine-tune is reduced, TTA still requires an additional backward propagation. Typically, the time cost of a backward propagation is approximately twice that of a forward propagation. Thus, Niu et al. (2024) propose FOA, which is free from backward propagation by adapting soft prompt through covariance matrix adaptation evolution strategy.
The stability of TTA is primarily shown in two aspects. On the one hand, unsupervised or self-supervised learning signals inevitably introduce noise into the optimization process, resulting in TTA optimizing the model in the incorrect gradient direction. To address this, Niu et al. (2023) and Gong et al. (2024a) propose noise data filtering strategies and the robust sharpness-aware optimizer. On the other hand, in real-world scenarios, the distribution of test samples may continually shift, but continual TTA optimization may lead to catastrophic forgetting of the model’s original knowledge. Episodic TTA (Shu et al., 2022; Zhao et al., 2024a) is a setting to avoid forgetting, which resets the model parameters to their original state after TTA on a single test sample. However, episodic TTA frequently loads the original model, leading to higher inference latency and also limiting the model’s incremental learning capability. To overcome the dilemma, a common trick is the exponential moving average (Wortsman et al., 2022; Ye et al., 2022), which incorporates information from previous model states.
3.2. Modifying the Input
When it comes to LLM, the large number of parameters makes model update-based TTA methods face a tougher dilemma of efficiency and stability. As a result, input-modification-based methods, which do not rely on parameter updates, have become the mainstream method for TTA in LLMs. The effectiveness of input-modified TTA stems from the in-context learning (ICL) capability of LLM, which can significantly improve the performance by adding some demonstrations before the test sample. ICL is highly sensitive to the selection and order of demonstrations. Therefore, the core objective of input-modification TTA is to select appropriate demonstrations for the test samples and arrange them in the optimal order to maximize the effectiveness of ICL.
First, empirical studies (Liu et al., 2022) show that the more similar the demonstrations are to the test sample, the better the ICL performance. Therefore, retrieval models like BM25 and SentenceBERT are used to retrieve demonstrations semantically closest to the test sample and rank them in descending order of similarity (Qin et al., 2024b; Luo et al., 2023b). To improve the accuracy of demonstration retrieval, Rubin et al. (2022) and Li et al. (2023b) specifically train the demonstration retriever by contrastive learning. Then, as researchers delve deeper into the mechanisms of ICL, ICL is considered to conduct implicit gradient descent on the demonstrations (Dai et al., 2023). Therefore, from the perspective of training data, demonstrations also need to be informative and diverse (Su et al., 2022; Li and Qiu, 2023). Wang et al. (2023d) view language models as topic models and formulate the demonstration selection problem as solving a Bayesian optimal classifier. Additionally, the ordering of examples is another important area for improvement. Lu et al. (2022) and Wu et al. (2023) use information theory as a guide to select the examples with maximum local entropy and minimum description length for ranking, respectively. Scarlatos and Lan (2024) and Zhang et al. (2022a) consider the sequential dependency among demonstrations, and model it as a sequential decision problem and optimize demonstration selection and ordering through reinforcement learning.
Another line of work (Chen et al., 2023; Lyu et al., 2023; Kim et al., 2022; Zhang et al., 2023d) argues that in practice, combining a limited set of externally provided examples may not always be the optimal choice. LLMs can leverage their generative and annotation capabilities to create better demonstrations. DAIL (Su et al., 2024) constructs a demonstration memory, storing previous test samples and their predictions as candidate demonstrations for subsequent samples. DAWN-ICL (Tang et al., 2024) further models the traversal order of test samples as a planning task and optimizes it by the Monte Carlo tree search (MCTS).
3.3. Editing the Representation
For generative LLMs, some works have found that the performance bottleneck is not in encoding world knowledge, but in the large gap between the information in intermediate layers and the output. During the inference phase, editing the representation can help externalize the intermediate knowledge into the output. PPLM (Dathathri et al., 2020) performs gradient-based representation editing under the guidance of a small language model to control the style of outputs. ActAdd (Turner et al., 2024) selects two semantically contrastive prompts and calculates the difference between their representations as a steering vector, which is then added to the residual stream. Representation editing based on contrastive prompts has demonstrated its effectiveness in broader scenarios, including instruction following (Stolfo et al., 2024), alleviating hallucinations (Li et al., 2023c; Arditi et al., 2024), reducing toxicity (Liu et al., 2024e; Lu and Rimsky, 2024) and personality (Cao et al., 2024). SEA (Qiu et al., 2024b) projects representations onto directions with maximum covariance with positive prompts and minimum covariance with negative prompts. They also introduce nonlinear feature transformations, allowing representation editing to go beyond linearly separable representations. Scalena et al. (2024) conduct an in-depth study on the selection of steering intensity. They find that applying a gradually decreasing steering intensity to each output token can improve control over the generation without compromising quality.
3.4. Calibrating the Output
Using external information to calibrate the model’s output distribution is also an efficient yet effective test-time adaptation method (Khandelwal et al., 2020). AdaNPC (Zhang et al., 2023c) designs a memory pool to store training data. During inference, given a test sample, AdaNPC recalls $k$ samples from the memory pool and uses a $k$NN classifier to predict the test sample. It then stores the test sample and its predicted label in the memory pool. Over time, the sample distribution in the memory pool gradually aligns with the test distribution. In NLP, the most representative application of such methods is $k$NN machine translation ($k$NN-MT). $k$NN-MT (Khandelwal et al., 2021) constructs a datastore to store contextual representations and their corresponding target tokens. During translation inference, it retrieves the $k$-nearest candidate tokens from the datastore based on the decoded context and processes them into probabilities. Finally, it calibrates the translation model’s probability distribution by performing a weighted fusion of the model’s probabilities and the retrieved probabilities. $k$NN-MT has demonstrated superior transferability and generalization compared to traditional models in cross-domain and multilingual MT tasks. Subsequent studies have focused on improving its performance and efficiency (Wang et al., 2022a; Zhu et al., 2023a; You et al., 2024) or applying its methods to other NLP tasks (Wang et al., 2022b; Bhardwaj et al., 2023).
Summary 1:
Parameter updating and output calibration are the most versatile TTA methods. However, parameter updating suffers from training instability and inefficiency in LLMs, while output calibration relies on target domain information and risks knowledge leakage. Input modification and representation editing are free from training but have limited applicability: input modification is related to ICL capabilities, and representation editing demands manual prior knowledge.
4. Test-time Reasoning for System-2 Thinking
Figure 4. Illustration of feedback modeling, search strategies and improvement training in test-time reasoning.
Test-time reasoning aims to spend more inference time to search for the most human-like reasoning process within the vast decoding search space. In this section, we introduce the two core components of test-time reasoning: feedback modeling and search strategies (as shown in Figure 4).
4.1. Feedback Modeling
Score-based Feedback
Score-based feedback, also known as the verifier, aims to predict scalar scores to evaluate the alignment of generated results with ground truth or human cognitive processes. Its training process is typically similar to the reward model in RLHF, using various forms of feedback signals and modeling it as a classification (Cobbe et al., 2021) or rank task (Bradley and Terry, 1952; Yuan et al., 2024a; Hosseini et al., 2024). In reasoning tasks, verifiers are primarily categorized into two types: outcome-based verifiers (ORMs) and process-based verifiers (PRMs). ORMs (Cobbe et al., 2021) use the correctness of the final CoT result as training signals. Liu et al. (2025b) provide a detailed recipe for training a strong ORM. In contrast, PRMs (Uesato et al., 2022; Lightman et al., 2024; Zhang et al., 2024f) are trained based on the correctness of each reasoning step. Compared to ORMs, PRMs cannot only evaluate intermediate reasoning steps but also assess the entire reasoning process more precisely. However, PRMs require more human effort to annotate feedback for the intermediate steps. Math-Shepherd (Wang et al., 2024g), OmegaPRM (Luo et al., 2024) and EpicPRM (Sun et al., 2025) utilize MCTS algorithm to collect high-quality process supervision data automatically. Zhang et al. (2025c) utilize critic models to evaluate process annotations collected by MCTS, filtering out low-quality data to improve the training effectiveness of PRMs. Setlur et al. (2024) argue that PRMs should evaluate the advantage of each step for subsequent reasoning rather than focusing solely on its correctness. They propose process advantage verifiers (PAVs) and efficiently construct training data through MCTS. Furthermore, Lu et al. (2024) and Yuan et al. (2024b) notice that ORMs implicitly model the advantage of each step, leading them to automatically annotate process supervision data using ORMs or directly train PRMs on outcome labels, respectively. Liu et al. (2025c) adaptively segment reasoning steps based on tokens with higher uncertainty, which is used to train better PRMs.
Generative-based Feedback
Although the verifier can evaluate the correctness of generated answers or steps, it lacks interpretability, making it unable to locate the specific cause of errors or provide correction suggestions. Generative-based feedback, also referred to critic, fully leverages the LLM’s generative and instruction-following ability (Ankner et al., 2024; Ye et al., 2024; Chiang et al., 2025; She et al., 2025). By designing specific instructions, it can perform pointwise or pairwise evaluation from multiple dimensions, and even provide suggestions for revision in natural language. Powerful closed-source LLMs, such as GPT-4 and Claude, are effective critics. They can perform detailed and controlled assessments of generated texts, such as factuality, logical errors, coherence, and alignment, with high consistency with human evaluations (Wang et al., 2023a; Luo et al., 2023a; Liu et al., 2023; Chiang and Lee, 2023). However, they still face biases such as length, position, and perplexity (Bavaresco et al., 2024; Wang et al., 2024f; Stureborg et al., 2024). LLM-as-a-Judge (Zheng et al., 2023) carefully designs system instructions to mitigate the interference of biases. BSM (Saha et al., 2024) evaluates based on multiple criteria and then merges them. Peng et al. (2025a) employ multi-agents to jointly evaluate answers’ factuality and instruction-following.
To obtain cheaper verbal-based feedback, open-source LLMs can also serve as competitive alternatives through supervised fine-tuning  (SFT) (Wang et al., 2024n; Zhu et al., 2023b; Liang et al., 2024b; Paul et al., 2024). Shepherd (Wang et al., 2023c) collects high-quality training data from human annotation and online communities to fine-tune an evaluation model. Auto-J (Li et al., 2024f) collects queries and responses from various scenarios and designs evaluation criteria for each scenario. GPT-4 then generates critiques of the responses based on these criteria and distills its critique ability to open-source LLMs. Prometheus (Kim et al., 2024c, d) designs more fine-grained evaluation dimensions. It trains a single evaluation model and a pairwise ranking model separately, then unifies them into one LLM by weight merging. To reduce reliance on human annotations and external LLMs, Wang et al. (2024d) propose a self-training method: the critique model generates positive and negative responses, then collects critique data via rejection sampling to perform iterative finetuning. Building on self-training, EvalPlanner enables (Saha et al., 2025) the critique model to plan evaluation processes and criteria, conduct critiques based on these, and then collect positive and negative samples to improve the critique model via DPO (Rafailov et al., 2023). (Yu et al., 2025b) carefully synthesize and filter data, enabling the base model to achieve strong critique abilities with only 40K samples for SFT and DPO training. GenRM (Zhang et al., 2024c) leverages instruction tuning to enable the verifier to answer ‘Is the answer correct (Yes/No)?’ and uses the probability of generated ‘Yes’ token as the score. GenRM can also incorporate CoT, allowing the verifier to generate the corresponding rationale before answering ‘Yes’ or ‘No’. ThinkPRM (Khalifa et al., 2025) evaluates each reasoning step with long CoT and requires only 1% of the process supervision data compared to discriminative PRMs. Critic-RM (Yu et al., 2024a) jointly trains the critic and the verifier.
Category Sub-category Representative Methods Domain Objective Description
Score-based ORM Cobbe et al. (2021) Math classification ORM; human annotated data
Acemath (Liu et al., 2025b) Math list-wise Bradley-Terry ORM; sampling training data from multiple LLMs
PRM Lightman et al. (2024) Math classification PRM; human annotated data
Math-shepherd (Wang et al., 2024g) Math classification PRM; annotating processes via MC estimation
Zhang et al. (2025c) Math classification/regression PRM; annotating processes via MCTS and LLM-as-a-judge
Implicit PRM (Yuan et al., 2024b) Math implicit reward modeling PRM; training PRMs with outcome labels
ASPRM (Liu et al., 2025c) Math, Code classification PRM; adaptive segment step; annotating processes via MC estimation
Generative-based Training-free LLM-as-a-Judge (Zheng et al., 2023) General - Designing system instructions to mitigate biases
BSM (Saha et al., 2024) General - Dividing into multiple criteria and then merging
Training-based Shepherd (Wang et al., 2023c) General SFT Collecting data from human annotation and the Internet
Prometheus (Kim et al., 2024d) General SFT Training single and pairwise models and then merging them
EvalPlanner (Saha et al., 2025) General DPO Planing evaluation processes and then evaluating
GenRM (Zhang et al., 2024c) Math SFT PRM; synthesizing critique data from external LLMs
R-PRM (She et al., 2025) Math SFT & DPO PRM; synthesizing critique data from external LLMs
Critic-RM (Yu et al., 2024a) General SFT & Bradley-Terry ORM; synthesizing and filtering critique data via self-critique
CLoud (Ankner et al., 2024) General SFT & Bradley-Terry ORM; synthesizing data from external LLMs and self-critique
Table 1. Overview of feedback modeling methods.
Category sub-category Representative Methods Tasks Verifier/Critic Train-free
Repeat Sampling Majority voting CoT-SC (Wang et al., 2023b) Math, QA self-consistency $\checkmark$
PROVE (Toh et al., 2024) Math compiler $\checkmark$
Best-of-N Cobbe et al. (2021) Math ORM
DiVeRSe (Li et al., 2023a) Math PRM
Knockout (Liu et al., 2025f) Math critic $\checkmark$
Self-correction Human feedback NL-EDIT (Elgohary et al., 2021) Semantic parsing Human
FBNET (Tandon et al., 2022) Code Human
External tools DrRepair (Yasunaga and Liang, 2020) Code compiler
Self-debug (Chen et al., 2024c) Code compiler $\checkmark$
CRITIC (Gou et al., 2024) Math, QA, Detoxifying text-to-text APIs $\checkmark$
External models REFINER (Paul et al., 2024) Math, Reason critic model
Shepherd (Wang et al., 2023c) QA critic model
Multiagent Debate (Du et al., 2023b) Math, Reason multi-agent debate $\checkmark$
MAD (Liang et al., 2024a) Translation, Math multi-agent debate $\checkmark$
Intrinsic feedback Self-Refine (Madaan et al., 2023) Math, Code, Controlled generation self-critique $\checkmark$
Reflexion (Shinn et al., 2023) QA self-critique $\checkmark$
RCI (Kim et al., 2023) Code, QA self-critique $\checkmark$
Tree Search Uninformed search ToT (Yao et al., 2023) Planing, Creative writing self-critique $\checkmark$
Xie et al. (2023) Math self-critique $\checkmark$
Heuristic search RAP (Hao et al., 2023) Planing, Math, Logical self-critique $\checkmark$
TS-LLM (Feng et al., 2024a) Planing, Math, Logical ORM
rStar (Qi et al., 2024a) Math, QA multi-agent consistency $\checkmark$
ReST-MCTS* (Zhang et al., 2024m) Math, QA PRM
Table 2. Overview of search strategies.
4.2. Search Strategies
4.2.1. Repeated Sampling
Sampling strategies such as top-p and top-k are commonly used decoding algorithms in LLM inference. They introduce randomness during decoding to enhance text diversity, allowing for parallelly sampling multiple generated texts. Through repeated sampling, we have more opportunities to find the correct answer. Repeated sampling is particularly suitable for tasks that can be automatically verified, such as code generation, where we can easily identify the correct solution from multiple samples using unit tests (Li et al., 2022a; Rozière et al., 2024). For tasks that are difficult to verify, like math word problems, the key to the effectiveness of repeated sampling is the verification strategy.
Verification strategy
Verification strategies include two types: majority voting and best-of-N (BoN) sampling. Majority voting (Li et al., 2024h; Lin et al., 2024a) selects the most frequently occurring answer in the samples as the final answer, which is motivated by ensemble learning. Majority voting is simple yet effective. For instance, self-consistency CoT (Wang et al., 2023b) can improve accuracy by 18% over vanilla CoT in math reasoning tasks. However, the majority does not always hold the truth, as they may make similar mistakes. Therefore, some studies perform filtering or weighting before voting. For example, the PROVE framework (Toh et al., 2024) converts CoT into executable programs, filtering out samples if the program’s results are inconsistent with the reasoning chain’s outcomes. Kang et al. (2025b) propose the self-certainty metric to weight votes based on their ranking. Huang et al. (2025a) use the calibrated confidence as vote weights, and adjust the number of samples based on response agreement to improve efficiency (Aggarwal et al., 2023). RPC (Zhou et al., 2025) demonstrates theoretically and experimentally that filtering low-probability samples improves the performance and efficiency of majority voting.
Best-of-N sampling uses a verifier to score each response and selects the one with the highest score as the final answer (Stiennon et al., 2020; Cobbe et al., 2021; Nakano et al., 2022; Liu et al., 2025a). Li et al. (2023a) propose a voting-based BoN variant, which performs weighted voting on all answers based on the verifier’s scores and selects the answer with the highest score. (Liu et al., 2025f) design BoN in a knockout tournament, using pairwise comparison verifiers to filter out the best response. In addition, some works aim to improve the efficiency of BoN. Inspired by speculative decoding, Zhang et al. (2024b); Qiu et al. (2024a); Sun et al. (2024); Yu et al. (2025a) and Manvi et al. (2024) evaluate each reasoning step and prune low-scoring sampled results, halting further generation for those paths, thereby significantly reducing the overall time cost. PRS (Ye and Ng, 2024) enables LLMs to self-critique and self-correct, guiding the model to generate expected responses with fewer sampling times. Li et al. (2025a) compare BoN and majority voting, demonstrating that BoN is suitable for harder questions with moderate diversity of response distribution.
Improvement Training
Repeated sampling has proven to be a simple yet effective method, even surpassing models fine-tuned with RLHF (Gao et al., 2023b; Hou et al., 2024). However, it costs much inference time that is difficult to afford in practical applications. Therefore, many studies have attempted to train the model by BoN sampling to approximate the BoN distribution, thereby reducing the search space during inference. ReST (Gulcehre et al., 2023) samples responses with reward values above a threshold from the policy model as self-training data and fine-tunes the policy model by offline reinforcement learning. In each iteration, ReST samples new training data. vBoN (Amini et al., 2024), BoNBoN (Gui et al., 2024) and BOND (Sessa et al., 2024) derive the BoN distribution and minimize the difference between the policy model’s distribution and the BoN distribution. Chow et al. (2024) design a BoN-aware loss to make the policy model more exploratory during fine-tuning.
4.2.2. Self-correction
Self-correction is a sequential test-time compute method that enables LLMs to iteratively revise and refine generated results based on external or internal feedback (Shinn et al., 2023).
Figure 5. Illustration of self-correction with tool checking, external model evaluation, and self-critique.
Feedback sources
The feedback used for self-correction is typically presented in natural language and comes from various sources, including human evaluation, tool checking, external model evaluation, and intrinsic feedback. Human evaluation is the gold standard for feedback, but due to its high cost and limited scalability, it is mainly used in early research to explore the upper limits of self-correction capabilities (Tandon et al., 2021; Elgohary et al., 2021; Tandon et al., 2022). For certain domain-specific tasks, external tool checking provides accurate and efficient feedback (Gou et al., 2024; Chen et al., 2024c; Gao et al., 2023a). For example, Yasunaga and Liang (2020) propose to obtain feedback from compilers in code repair and generation tasks. In embodied tasks, the environment can provide precise feedback on the action trajectories of LLM-based agents (Wang et al., 2024e).
External model evaluation is an effective feedback source for general tasks, such as various verbal-based critique models described in Section 4.1. For example, Paul et al. (2024) first define multiple error types for natural language reasoning tasks and then design the corresponding feedback templates. They train an evaluation model using synthetic feedback training data, and with the critic, the reasoning model achieves substantial performance improvement. Multi-agent debate (Du et al., 2023b; Xiong et al., 2023; Liang et al., 2024a; Chen et al., 2024e; Wang et al., 2024l) is another mechanism that leverages external feedback to enhance reasoning capabilities. In this approach, models do not have distinct roles as reasoners and critics. Instead, multiple models independently conduct reasoning, critique each other, and defend or refine their reasoning based on feedback. This process continues until agents reach a consensus or a judge model summarizes the final reasoning results. The multi-agent debate has shown its potential in fact-checking (Kim et al., 2024b; Khan et al., 2024), commonsense QA (Xiong et al., 2023), faithful evaluations (Chan et al., 2024), and complex reasoning (Du et al., 2023b). However, multi-agent debate may be unstable, as LLMs are susceptible to adversarial information and may revise correct answers to incorrect ones in response to misleading inputs (Laban et al., 2024; Amayuelas et al., 2024). Therefore, a successful multi-agent debate requires that LLMs maintain their stance when faced with incorrect answers from other models while remaining open to valid suggestions (Stengel-Eskin et al., 2024). In general, the more LLMs involved in the debate, the stronger the overall reasoning performance. However, this significantly increases the number of LLM inferences required, and the length of input context, posing a major challenge to LLM inference costs (Liu et al., 2024b). To reduce debate inference costs, Li et al. (2024b) investigate the impact of topological connections among multiple agents and show that sparse connections, such as ring structures, are not inferior to the fully connected topology. GroupDebate (Liu et al., 2024b) divides LLMs into groups that conduct debates internally and only share the consensus results between groups.
Self-critique assumes that LLMs can self-evaluate their outputs and optimize them through intrinsic feedback (Yuan et al., 2024c). This idea stems from a fundamental principle in computational complexity theory: verifying whether a solution is correct is typically easier than solving the problem. Bai et al. (2022) propose self-correcting harmful responses from LLMs by prompting themselves. Self-Refine (Madaan et al., 2023) and RCI Prompting (Kim et al., 2023) iteratively prompt LLMs to self-correct their responses in tasks such as arithmetic reasoning. IoE (Li et al., 2024a) observes that LLMs may over-criticize themselves during self-critique, leading to performance degradation, and designs prompt to guide LLMs in assessing confidence. ProgCo (Song et al., 2025b) leverage the advantages of code in expressing complex logic, enabling LLMs to generate responses in pseudo-code form, followed by self-critique and refinement. SETS (Chen et al., 2025c) combines the strengths of repeated sampling and self-critique, applying self-critique and correction to each sampled reasoning path and choosing the final solution via majority voting. S1 (Muennighoff et al., 2025) adds the “wait” to the reasoning process, prompting the LLM to critique its reasoning.
Arguments
The effectiveness of self-correction, especially the self-critique, has remained controversial. Several empirical studies on code generation (Olausson et al., 2024), commonsense QA (Huang et al., 2024a), math problem-solving (Wang et al., 2024c), planning (Valmeekam et al., 2023a), and graph coloring (Stechly et al., 2023) confirm that self-correction is not a guaranteed solution for improving performance. Kamoi et al. (2024) think the effectiveness of self-correction has been overestimated. Previous successes either rely on oracle answers or weak initial answers. Only tasks that can be broken down into easily verifiable sub-tasks can truly benefit from self-correction. They suggest fine-tuning specific evaluation models to achieve better self-correction. Zhang et al. (2024e) try to interpret and alleviate the failure of self-critique via human-like cognitive bias. Tyen et al. (2024) decouple the abilities of LLMs to identify and correct errors and create the corresponding evaluation datasets. The evaluation results show that LLMs do not lack the ability to correct errors during self-correction, and their main performance bottleneck lies in locating the errors. Yang et al. (2024b) decompose self-critique into confidence and critique capabilities. Empirical studies show that fine-tuning is necessary to enhance both capabilities simultaneously, while prompt engineering can only achieve a trade-off.
Improvement Training
Most of the aforementioned self-correction methods demonstrate significant performance improvements on advanced closed-source large models or open-source LLMs with over 70B parameters. However, for medium-scale open-source models with weaker capabilities, we need to further fine-tune them to unlock their self-correction capabilities. Supervised fine-tuning optimizes the model using high-quality multi-turn correction data, either manually annotated (Saunders et al., 2022), self-rationalize (Zelikman et al., 2022; Yuan et al., 2025a), multi-agent debate (Subramaniam et al., 2025) or sampled from stronger LLMs (An et al., 2023; Paul et al., 2024; Qu et al., 2024; Gao et al., 2024a; Zhang et al., 2024a; Xi et al., 2024). GLoRe (Havrilla et al., 2024) considers that LLMs need global or local refinement for different types of errors. To address this, they construct training sets for global and local refinement, train verifiers to identify global and local errors, and develop LLMs for refinement based on different global or local feedback signals. Xi et al. (2024) design a scalable framework for synthesizing self-correction training data, enabling reasoning models to generate controlled errors and receive feedback from critics to self-correct. Although SFT is effective, training data from offline-generated self-correction trajectories can only simulate limited correction patterns. This leads to the distribution mismatch with the actual self-correction behavior during model inference. Self-correct (Welleck et al., 2023) adopts online imitation learning, re-sampling new self-correction trajectories for training after each training epoch. To further expand the exploration space of LLMs, many studies adopt flexible RL algorithms to surpass the performance limits of SFT. SCoRe (Kumar et al., 2024) proposes using the multi-turn RL method to improve self-critique and self-correction capability. T1 (Hou et al., 2025) employs self-correction training data for SFT cold-start, followed by RL training using the RLOO algorithm (Ahmadian et al., 2024). During the RL phase, high-temperature sampling and entropy rewards encourage the LLM to explore more diverse reasoning paths. Deepseek-R1 (Guo et al., 2025a) uses rule-based rewards and the GRPO algorithm (Shao et al., 2024) for RL training. It also demonstrates RL’s immense potential, even without SFT cold-start, its exploration capabilities suffice to endow LLMs with strong reasoning abilities.
4.2.3. Tree Searching
Repeated sampling and self-correction scale test-time compute in parallel and sequentially, respectively. Human thinking is a tree search that combines brainstorming in parallel with backtracking to find other paths to solutions when it encounters a dead end. Search algorithms and value functions are two critical components in tree searching.
Search algorithm
In LLM reasoning, current search algorithms include uninformed search and heuristic search. Uninformed search explores the search space according to a fixed rule. For example, tree-of-thought (ToT) (Yao et al., 2023) adopts the BFS or DFS to search, while Xie et al. (2023) use beam search. Uninformed search is usually less efficient for problems with large search spaces, so heuristic search strategies represented by A (Meng et al., 2024; Wang et al., 2024a) and MCTS (Hao et al., 2023; Bi et al., 2024; Park et al., 2024) are widely used in reasoning tasks. MCTS, which eliminates the need for explicit heuristics, leverages stochastic simulations and adaptive tree expansion under uncertain environments, making it well-suited for large state spaces. It optimizes search results gradually through four steps: selection, expansion, simulation, and backpropagation, approaching the optimal solution. In contrast, A uses a heuristic function-guided deterministic search to guarantee optimal paths, but its performance depends on the design of the heuristic function. As a result, MCTS has been successfully applied to tasks such as RAG (Hu et al., 2024b; Jiang et al., 2024; Li and Ng, 2024; Feng et al., 2025), QA (Luo et al., 2025a; Gan et al., 2025), hallucinations mitigation (Cheng et al., 2025), text-to-SQL (Yuan et al., 2025b), etc. Additionally, Long (2023) trains an LLM controller using reinforcement learning to guide the LLM reasoner’s search path, and Chari et al. (2025) utilizes ant colony evolutionary algorithm to guide tree search.
Value function
The value function evaluates the value of each state and guides the tree to expand towards branches with higher values in heuristic tree search. Xu (2023) train an energy function by noise-contrastive estimation as the value function. RAP (Hao et al., 2023) designs a series of heuristic value functions, including the likelihood of the action, the confidence of the state, self-evaluation results, and task-specific reward, and combines them according to task requirements. Reliable and generalized value functions facilitate the application of MCTS to more complex problems with deeper search spaces. AlphaMath (Chen et al., 2024b) and TS-LLM (Feng et al., 2024a) replace the hand-crafted value function with a learned LLM value function, automatically generating reasoning process and step-level evaluation signals in MCTS. VerifierQ (Qi et al., 2024b) integrates implicit Q-learning and contrastive Q-learning to train the value function, effectively mitigating the overestimation issue at the step level. Traditional MCTS methods expand only one trajectory, while rStar (Qi et al., 2024a) argues that the current value function struggles to guide the selection of the optimal path accurately. Therefore, rStar retains multiple candidate paths and performs reasoning with another LLM, ultimately selecting the path where both LLMs’ reasoning results are consistent. Gao et al. (2024c) propose SC-MCTS, which combines multiple reward models, including contrastive reward, likelihood, and self-evaluation as value functions. MCTSr (Zhang et al., 2024d) and SR-MCTS (Zhang et al., 2024g) take complete responses as nodes, expanding the search space through self-critique and correction. SR-MCTS utilizes pairwise preference rewards and global quantile score as the value function, offering a more robust value function estimation.
Improvement Training
Tree search can guide LLMs to generate long reasoning processes, and these data help train LLMs with stronger reasoning abilities (Zhai et al., 2024; Xu et al., 2024b; Guan et al., 2025). ReST-MCTS* (Zhang et al., 2024m) uses process rewards as a value function to guide MCTS, collecting high-quality reasoning trajectories and the value of each step to improve the policy model and reward model. Due to the step-by-step exploration of tree search, it can obtain finer-grained step-level feedback signals. MCTS-DPO (Xie et al., 2024) collects step-level preference data through MCTS and uses DPO for preference learning. AlphaLLM-CPL (Wang et al., 2024k) ranks trajectories based on preference reward gaps and policy prediction gaps, employing curriculum learning to efficiently utilize MCTS-collected trajectories. Recently, many LMRs (Qin et al., 2024a; Zhao et al., 2024b; Zhang et al., 2024h) have also confirmed the necessity of using tree search to construct high-quality long reasoning chain data for training.
Summary 2:
Repeated sampling is easy to implement and improves answer diversity, making it suitable for open-ended or easily verifiable tasks, though computationally inefficient. Self-correction relies on precise, fine-grained feedback and works well for easily verifiable tasks, but may not perform well with poor feedback or weak reasoning capability. Tree search optimizes complex planning tasks globally but involves complex implementation.
5. Advanced Topics
5.1. Generalizable System-2 Model
Currently, most LRMs exhibit strong reasoning performance in specific domains such as math and code, but they struggle to generalize to cross-domain, cross-lingual, or general tasks (Son et al., 2025; Zhao et al., 2025b). On the one hand, as the foundation of current System-2 models, CoT shows little effectiveness in non-symbolic reasoning tasks (Sprague et al., 2024). On the other hand, verifiers and critics have limited generalization capabilities, making it difficult to provide effective guidance to the reasoning model (LeVine et al., 2023; Kim et al., 2024a; Chen et al., 2024f). For the former limitation, integrating external tools and multi-agent feedback is a promising direction (Nathani et al., 2023; Lan et al., 2024). Many works aim to enable reasoning models to learn to perform retrieval or tool use during the reasoning process. Search-o1 (Li et al., 2025b) triggers external search tools by generating special tokens that represent search actions, and Search-R1 (Jin et al., 2025) further enhances the interaction between the LLM and search tools through RL. Deep Research (OpenAI, 2025; Google, 2024; Zheng et al., 2025; Li et al., 2025e) integrates various tools such as web search and code execution, demonstrating distinguished performance on general tasks like information synthesis, report writing, and complex code generation. For the latter challenge, some works utilize multi-objective training (Wang et al., 2024m), model ensemble (Lin et al., 2024c), soft reward (Su et al., 2025) or regularization constraints (Yang et al., 2024a; Jia, 2024) to make verifiers more generalizable. Compared to discriminative verifiers, generative critics inherit the generalization ability of LLMs and can leverage test-time compute to realize greater potential (Kim et al., 2025). GenPRM (Zhao et al., 2025a) and GRM (Liu et al., 2025e) leverage repeated sampling for test-time compute, and GRM trains a Meta RM to evaluate the quality of the critic’s outputs. RM-R1 (Chen et al., 2025b) and JudgeLRM (Chen et al., 2025a) use RL training to enable the critic with self-correction ability.
Additionally, weak-to-strong generalization (Burns et al., 2023) is a topic worth further exploration. People are no longer satisfied with solving mathematical problems with standard answers; they hope System-2 models can assist in scientific discovery and the proofs of mathematical conjectures. In such cases, even human experts struggle to provide accurate feedback, while weak-to-strong generalization offers a promising direction to address this issue (Tang et al., 2025).
5.2. Multimodal Test-time Compute
Visual, audio, video and other modalities are crucial for models understanding and interaction with the world. Like LLMs, large multimodal models (LMMs) also face challenges in scaling training compute. Thus, enabling efficient test-time scaling for multimodal tasks has become an advanced research topic. In System-1 thinking, TTA has been successfully applied to LMMs, improving performance in tasks such as zero-shot image classification, image-text retrieval, and image captioning (Zhao et al., 2024a). For multimodal reasoning tasks, the exploration of multimodal CoT (Zhang et al., 2024j; Wu et al., 2024b; Mondal et al., 2024; Lee et al., 2024; Gao et al., 2024b; Zhang et al., 2024k) and multimodal critics or verifiers (Xiong et al., 2024; Tu et al., 2025) open up the possibility of building multimodal System-2 models. For example, IoT (Zhou et al., 2024b) and SketchPAD (Hu et al., 2024a) utilize visual tools to obtain critical visual information and integrate it into CoT to refine multimodal reasoning. VisualPRM (Wang et al., 2025a) constructs process annotation data through Monte Carlo estimation and trains a lightweight PRM verifier; and cutting-edge LMMs such as GPT-4o and Qwen-VL can serve as critics. Building on these foundations, repeated sampling (Lin et al., 2025) and tree search (Yao et al., 2024; Dong et al., 2024b) can effectively further improve LMMs’ multimodal reasoning performance. Xu et al. (2024a) divide the visual reasoning process into four stages: task summary, caption, reasoning, and answer conclusion. They propose a stage-level beam search method, which repeatedly samples at each stage and selects the best result for the next stage. Audio-Reasoner (Xie et al., 2025b) transfers the four-stage CoT framework to the audio modality. To equip LMMs with the self-correction ability, Deepseek-R1’s training recipe has also been practiced in LMMs (Ma et al., 2025a; Chen et al., 2025d; Shen et al., 2025a). OpenVLThinker (Deng et al., 2025) and Vision-R1 (Huang et al., 2025b) convert images into text captions, enabling text-only LRMs to incorporate visual information for generating CoTs. This serves as cold-start data for further reinforcement learning to train multimodal reasoning models. LMM-R1 (Peng et al., 2025b) demonstrates that the reasoning ability acquired through text-only reinforcement learning can serve as an effective cold start for multimodal reinforcement learning training.
Besides understanding and reasoning tasks, multimodal generation tasks can also benefit from test-time compute. Mainstream multimodal generative models are divided into diffusion-based (Peebles and Xie, 2023; Yang et al., 2025c) and autoregressive-based (Tian et al., 2024; Xie et al., 2025c) models. While diffusion models lack explicit CoTs, they can expand the search space at test time by sampling various Gaussian noises (Xie et al., 2025a). For video generation, due to the high computational cost of repeatedly sampling full videos, Cong et al. (2025) use PRM to verify frames individually and apply beam search to prune low-scoring frames. Liu et al. (2025d) propose frame tree search, using critics to evaluate videos at early, middle, and late stages with different criteria, retaining only the top-k scoring frames at each step. Autoregressive models treat intermediate generated images as reasoning steps. Guo et al. (2025b) propose a Potential Assessment Reward Model (PARM) to evaluate the quality potential of these intermediate images and introduce a reflection mechanism that enables the generative model to self-correct low-quality images.
5.3. Efficient Test-time Compute
The successful application of test-time compute shows that sacrificing reasoning efficiency can lead to better reasoning performance. However, researchers continue to seek a balance between performance and efficiency, aiming to achieve optimal performance under a fixed reasoning latency budget. This requires adaptively allocating computational resources for each sample. Damani et al. (2024) train a lightweight module to predict the difficulty of a question, and allocate computational resources according to its difficulty. Zhang et al. (2024l) further extend the allocation targets to more hyperparameters. Chen et al. (2025e) and Wang et al. (2025b) systematically evaluate the over-thinking and under-thinking phenomena in LRMs, where the former leads models to overcomplicate simple problems, and the latter causes frequent switching of reasoning paths on difficult problems, thereby reducing reasoning efficiency. These phenomena even cause LRM to perform poorly when dealing with simple problems (Zhang et al., 2025b). There are still many open questions worth exploring, such as how to integrate inference acceleration strategies, e.g. model compression (Li et al., 2024c; Huang et al., 2024c; Li et al., 2025f), token pruning (Fu et al., 2024; Zhang et al., 2024i), and speculative decoding (Leviathan et al., 2023; Xia et al., 2024) with test-time compute, and how to allocate optimal reasoning budget according to problem difficulty (Wang et al., 2024b; Han et al., 2024; Cheng and Durme, 2024).
Long CoT leads to inference latency and memory footprints of key-value cache. Recently, numerous studies have explored various strategies to reduce the reasoning length. Early work primarily focus on SFT models (Yu et al., 2024c; Liu et al., 2024a; Kang et al., 2025a; Xia et al., 2025; Ma et al., 2025c; Munkhbat et al., 2025; Cui et al., 2025; Yang et al., 2025a) using variable-length CoT data. In contrast, mainstream approaches incorporate length-based rewards into reinforcement learning (RL) (Team et al., 2025; Luo et al., 2025c; Arora and Zanette, 2025; Yeo et al., 2025; Yi et al., 2025; Yuan et al., 2025c) to encourage concise responses. Xiao et al. (2025) measure the complexity of vision-language reasoning tasks and design a complexity-aware curriculum learning and GRPO algorithm. They organize training samples from hard to easy and dynamically adjust the reward function and KL term weights based on task complexity. L1 (Aggarwal and Welleck, 2025) and Elastic Reasoning (Xu et al., 2025a) can enable precise control over response length based on a given token budget. Further, SelfBudgeter (Li et al., 2025c) can autonomously estimate the optimal token budget before generation, and allow users to choose whether to wait for the full response or terminate early during the generation process. However, both SFT and RL require substantial computational resources. Several studies have instead explored training-free approaches for efficient reasoning, including prompting (Han et al., 2024; Xu et al., 2025d; Aytes et al., 2025; Lee et al., 2025), model merging (Wu et al., 2025; Team et al., 2025) and early exit (Fu et al., 2025; Ma et al., 2025b; Yang et al., 2025b). Notably, DEER (Yang et al., 2025b) makes early-exit decisions during CoT generation based on the confidence of intermediate answers, effectively reducing reasoning overhead. Subsequently,  Dai et al. (2025) introduces a new RL paradigm that encourages the model to incentivize early thinking termination when appropriate. Another promising approach to achieving efficiency is hybrid reasoning, which dynamically switches between concise responses and long-chain reasoning based on task complexity. These methods (Lou et al., 2025; Luo et al., 2025b; Jiang et al., 2025; Fang et al., 2025; Zhang et al., 2025a) follow a similar pipeline: they start with SFT on a mixed dataset of long and short CoT samples as a cold start, and then refine the policy optimization algorithm to train the model to acquire hybrid reasoning capabilities.
6. Future Directions
6.1. Test-time Scaling Law
Unlike training-time computation scaling, test-time compute still lacks a universal scaling law. Some works have attempted to derive scaling laws for specific test-time compute strategies (Wu et al., 2024c; Levi, 2024). Brown et al. (2024) demonstrate that the performance has an approximately log-linear relationship with repeated sampling times. Chen et al. (2024d) models repeated sampling as a knockout tournament and league-style algorithm, proving theoretically that the failure probability of repeated sampling follows a power-law scaling. Snell et al. (2024) investigate the scaling laws of repeated sampling and self-correction, and propose the computing-optimal scaling strategy. There are two major challenges to achieving a universal scaling law: first, current test-time compute strategies are various, each with different mechanisms to steer the model; thus, it lacks a universal framework for describing them; second, the performance of test-time compute is affected by a variety of factors, including the difficulty of samples, the accuracy of feedback signals, and decoding hyperparameters, and we need empirical studies to filter out the critical factors.
6.2. Strategy Combination
Different test-time compute strategies are suited to various tasks and scenarios, so combining multiple strategies is one way to achieve better System-2 thinking. For example, Marco-o1 (Zhao et al., 2024b) combines the MCTS and self-correction, using MCTS to plan reasoning processes, and self-correction to improve the accuracy of each step. TPO (Li et al., 2025d) combines BoN sampling and self-correction. Moreover, test-time adaptation strategies in System-1 models can also be combined with test-time reasoning strategies. Akyürek et al. (2024) combine test-time training with repeated sampling. They further optimize the language modeling loss on test samples, then generate multiple candidate answers through data augmentation, and finally determine the answer by majority voting. They demonstrate the potential of test-time training in reasoning tasks, surpassing the human average on the ARC challenge. Therefore, we think that for LLM reasoning, it is crucial to focus not only on emerging test-time strategies but also on test-time adaptation methods. By effectively combining these strategies, we can develop System-2 models that achieve or surpass o1-level performance.
6.3. New Test-time Compute Paradigms
Latent CoT reasoning has emerged as a promising paradigm for test-time compute. By performing internal reasoning in latent space, it overcomes the limitations of conventional explicit CoT in terms of computational efficiency and abstract reasoning capability (Goyal et al., 2024; Xu et al., 2025b; Shen et al., 2025b; Xu et al., 2025c). Coconut (Hao et al., 2024) performs reasoning in a continuous latent space by directly feeding the model’s final hidden state as the next-step input embedding, thereby giving rise to advanced reasoning patterns. CCoT (Cheng and Durme, 2024) introduces a framework that generates continuous and variable-length contemplation tokens as compressed representations of reasoning chains, enabling efficient and accurate reasoning. LightThinker (Zhang et al., 2025d) improves LLM efficiency by compressing intermediate reasoning steps into compact gist tokens via specialized data construction and attention mask design, enabling the model to discard verbose chains while maintaining the ability to reason over condensed representations. Latent CoT currently faces several challenges, including the difficulty of direct supervision, limited generalization, and concerns over interpretability (Chen et al., 2025f). This calls for further exploration of token-level strategies and internal mechanisms to enable more robust and transparent latent reasoning.
7. Benchmarks
Test-time Adaptation
In System-1 models, distribution shifts include adversarial robustness, cross-domain and cross-lingual scenarios. In the field of CV, ImageNet-C (Hendrycks and Dietterich, 2019), ImageNet-R (Hendrycks et al., 2021a), ImageNet-Sketch (Wang et al., 2019) are common datasets for TTA. Yu et al. (2023) propose a benchmark to conduct a unified evaluation of TTA methods across different TTA settings and backbones on 5 image classification datasets. For NLP tasks, TTA is primarily applied in QA and machine translation tasks, with commonly used datasets such as MLQA (Lewis et al., 2020), XQuAD (Artetxe et al., 2020), MRQA (Fisch et al., 2019) and CCMatrix (Schwenk et al., 2021).
Feedback Modeling
RewardBench (Lambert et al., 2024) collects 20.2k prompt-choice-rejection triplets covering tasks such as dialogue, reasoning, and safety. It evaluates the accuracy of reward models in distinguishing between chosen and rejected responses. RM-Bench (Liu et al., 2024d) further evaluates the impact of response style on reward models. RMB (Zhou et al., 2024a) extends the evaluation to the more practical BoN setting, where reward models are required to select the best response from multiple candidates. CriticBench (Lin et al., 2024b) is specifically designed to evaluate a critic model’s generation, critique, and correction capabilities. For PRM, Song et al. (2025a) propose PRMBench, which evaluates PRMs whether they can identify the earliest incorrect reasoning step in math tasks. ProcessBench (Zheng et al., 2024) provides a more fine-grained evaluation, including redundancy, soundness, and sensitivity. In addition, there are benchmarks for evaluating multimodal feedback modeling, such as VL-RewardBench (Li et al., 2024g), VCR-Bench (Qi et al., 2025) and MJ-Bench (Chen et al., 2024a).
Test-time Reasoning
Reasoning capability is the core of System-2 models, including mathematics, code, commonsense, planning, etc (Zeng et al., 2024). Math reasoning is one of the most compelling reasoning tasks. With the advancements in LLM and test-time compute, the accuracy on some previously challenging benchmarks, like GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b), have surpassed the 90% mark. Thus, more difficult college admissions exam (Zhang et al., 2023b; Arora et al., 2023; Azerbayev et al., 2024) and competition-level (Gao et al., 2024d) math benchmarks have been proposed. Some competition-level benchmarks are not limited to textual modalities in algebra, logic reasoning, and word problems. For instance, OlympiadBench (He et al., 2024), OlympicArena (Huang et al., 2024b) and AIME (Zamil and Rabby, 2024) provide images for geometry problems, incorporating visual information to aid in problem-solving, while AlphaGeometry (Trinh et al., 2024) employs symbolic rules for geometric proofs. The most challenging benchmark currently is FrontierMath (Glazer et al., 2024), with problems crafted by mathematicians and covering major branches of modern mathematics. Even the most advanced o3 has not achieved 30% accuracy.
Code ability is a key aspect of LLM reasoning, with high practical value, covering code completion (Ding et al., 2023; Zhang et al., 2023a; Gong et al., 2024b), code reasoning (Gu et al., 2024; Wei et al., 2025), and code generation (Chen et al., 2021; Austin et al., 2021) tasks. Among these, code generation gains more attention. HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) provide natural language descriptions of programming problems, requiring LLMs to generate corresponding Python code and use unit tests for evaluation. MultiPL-E (Cassano et al., 2022) extend them to 18 program languages. EvalPlus (Liu et al., 2024c) automatically augments test cases to assess the robustness of the generated code. Recently, some studies collect benchmarks from open-source projects, which are closed to realistic applications and more challenging due to complex function calls, such as DS-1000 (Lai et al., 2023), CoderEval (Yu et al., 2024b), EvoCodeBench (Li et al., 2024d) and BigCodeBench (Zhuo et al., 2025).
Commonsense reasoning requires LLMs to possess both commonsense and reasoning abilities. StrategyQA (Geva et al., 2021) collects complex and subtle multi-hop reasoning questions. MMLU (Hendrycks et al., 2021b) and MMLU-Pro (Wang et al., 2024j) cover commonsense reasoning questions across various domains, including STEM, the humanities, the social sciences, etc. Planning aims to enable LLMs to take optimal actions based on the current state and environment to complete tasks. Current planning benchmarks primarily focus on synthetic tasks, such as Blocksworld (Valmeekam et al., 2023b), Crosswords, and Game-of-24 (Yao et al., 2023).
8. Conclusion
In this paper, we conduct a comprehensive survey of existing works on test-time compute. We introduce various test-time compute methods in System-1 and System-2 models, and look forward to future directions for this field. We believe test-time compute can help models handle complex real-world distributions and tasks better, making it a promising path for advancing LLMs toward cognitive intelligence. We hope this paper will promote further research.